OpenStudy (calculusxy):

Ella spent her whole allowance of $2.00, plus the 16¢ she had left over from last week, on bubble gum. If the pieces of gum had been a penny cheaper, she would have received three more pieces than she did. How many pieces did she actually buy? @satellite73

OpenStudy (anonymous):

i must be tired lets put \(x\) as the cost of a piece of gum so \(x-1\) is the cost if it is a penny cheaper then i think we have to solve \[\frac{2.16}{x}=\frac{2.16}{x-1}+3\] for \(x\)

OpenStudy (anonymous):

no that cannot be right

OpenStudy (calculusxy):

How do I solve for x?

OpenStudy (anonymous):

i think i have the 3 on the wrong side

OpenStudy (anonymous):

damn even that is wrong

OpenStudy (anonymous):

trial an error strikes again!

OpenStudy (anonymous):

\[216=2^3\times 3^3\] is the key

OpenStudy (anonymous):

try \(.09\) and \(.08\) and you will see that it works

OpenStudy (anonymous):

\[\frac{216}{8}=27\] and \[\frac{216}{9}=24\]